OpenStudy (anonymous):

Which equation shows the substitution method being used to solve the system of linear equations? x + y = 6 x = y + 2

pooja195 (pooja195):

Hi :) So we have 2 equations x + y = 6 <---equation 1 x = y + 2 <---equation 2 We know what x equals so we plug equation 2 into equation 1 \[\huge~\rm~y+2+y=6\]Can you solve for y? :)

OpenStudy (anonymous):

2

pooja195 (pooja195):

Good well done :) now we know the value of y all we need to do is plug that back into equation 2 x = 2+ 2

OpenStudy (anonymous):

so a/b/c/d A. x = (x – 6) + 2 B. x + (y + 2) = 6 C. (y + 2) + y = 6 D. x + y = y + 2

pooja195 (pooja195):

Oh in that case just look back at the steps i showed you

OpenStudy (anonymous):

so c

pooja195 (pooja195):

Correct :)

OpenStudy (anonymous):

yay thx

pooja195 (pooja195):

yw